# Fully Dynamic Matching and Ordered Ruzsa-Szemerédi Graphs

### v2

@@ Abstract
We study the fully dynamic maximum matching problem. In this problem, the goal is to efficiently maintain an approximate maximum matching of a graph that is subject to edge insertions and deletions. Our focus is particularly on algorithms that maintain the edges of a $(1-\varepsilon)$-approximate maximum matching for an arbitrarily small constant $\varepsilon>0$. Until recently, the fastest known algorithm for this problem required $\Theta(n)$ time per update where $n$ is the number of vertices. This bound was slightly improved to $n/(\log^{*}n)^{\Omega(1)}$ by Assadi, Behnezhad, Khanna, and Li [STOC’23] and very recently to $n/2^{\Omega(\sqrt{\log n})}$ by Liu [ArXiv’24]. Whether this can be improved to $n^{1-\Omega(1)}$ remains a major open problem.
In this paper, we present a new algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching. The update-time of our algorithm is parametrized based on the density of a certain class of graphs that we call Ordered Ruzsa-Szemerédi (ORS) graphs, a generalization of the well-known Ruzsa-Szemerédi graphs. While determining the density of ORS (or RS) remains a hard problem in combinatorics, we prove that if the existing constructions of ORS graphs are optimal, then our algorithm runs in $n^{1/2+O(\varepsilon)}$ time for any fixed $\varepsilon>0$ which would be significantly faster than existing near-linear in $n$ time algorithms.
@@ 1 Introduction
The dynamic matching problem has received a lot of attention over the last two decades [33, 4, 5, 32, 27, 18, 11, 12, 17, 16, 35, 21, 1, 13, 9, 10, 37, 14, 15, 34, 29, 26, 8, 20, 7, 19]. There is a relatively simple algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching, for any fixed $\varepsilon>0$, in just $O(n)$ time per update (see e.g. [27]). The update-time can be significantly improved if we worsen the approximation. For instance, a 1/2-approximation can be maintained in $\operatorname{poly}\log n$ time per update [36, 4, 9], or an (almost) 2/3-approximation can be maintained in $O(\sqrt{n})$ time [12]. Nonetheless, when it comes to algorithms with approximation ratio better than $2/3$, the update-time stays close to $n$. A slightly sublinear algorithm was proposed by Assadi, Behnezhad, Khanna, and Li [3] which runs in $n/(\log^{*}n)^{\Omega(1)}$ time per update and maintains a $(1-o(1))$-approximation. In a very recent paper, Liu [30] improved this to $n/2^{\Omega(\sqrt{\log n})}$ via a nice connection to algorithms for the online matrix-vector multiplication (OMv) problem. Despite this progress, the following remains a major open problem:
@@
Is it possible to maintain a $(1-\varepsilon)$-approximte maximum matching in a fully dynamic graph, for any fixed $\varepsilon>0$, in $n^{1-\Omega(1)}$ update-time?
We note that there is an orthogonal line of work on fully dynamic algorithms that instead of maintaining the edges of the matching, maintain only its size [7, 20, 19]. For this easier version of the problem, Bhattacharya, Kiss, and Saranurak [19] positively resolved the open problem above. However, their algorithm crucially relies on only estimating the size and does not work for the problem of maintaining the edges of the matching. We refer interested readers to [7] where the difference between the two versions of the problem is mentioned.
@@ 1.1 Our Contributions
Let $\varepsilon>0$ be fixed. There is a fully dynamic algorithm that maintains the edges of a $(1-\varepsilon)$-approximate maximum matching in $O\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}\operatorname{poly}(\log n)\right)$ amortized update-time. The algorithm is randomized but works against adaptive adversaries.
@@
The best lower bound on $\mathsf{RS}_{n}(cn)$— i.e., the densest known construction—is that of [22] from more than two decades ago which shows $\mathsf{RS}_{n}(cn)\geq n^{\Omega_{c}(1/\log\log n)}=n^{o(1)}$. This is indeed the densest known construction of ORS graphs we are aware of too. If this turns out to be the right bound, Theorem 1 implies a $(1-\varepsilon)$-approximation in $\widetilde{O}(n^{1/2+\varepsilon})$ time, an almost quadratic improvement over prior near-linear in $n$ algorithms of [3, 30]. In fact, we note that so long as $\mathsf{ORS}_{n}(cn)$ is moderately smaller than $n$ (say $\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))\ll n^{1-2\varepsilon}$) Theorem 1 implies a truly sublinear in $n$ update-time algorithm, positively resolving 1.
Finally, we note that there is a long body of work on proving conditional lower bounds for dynamic problems. For instance, the OMv conjecture can be used to prove that maintaining an exact maximum matching requires near-linear in $n$ update-time [28]. Adapting these lower bounds to the $(1-\varepsilon)$-approximate maximum matching problem has remained open since then. Our Theorem 1 implies that proving such lower bounds either requires a strong lower bound of near-linear on $\mathsf{ORS}_{n}(\varepsilon n)$, or requires a conjecture that implies this.
Contribution 2: Better Upper Bounds for ORS and RS:
Unfortunately there is a huge gap between existing lower and upper bounds for $\mathsf{RS}_{n}(cn)$ (and as a result also for $\mathsf{ORS}_{n}(cn)$). The best known upper bound on $\mathsf{RS}_{n}(cn)$ for linear size matchings follows from the improved triangle-removal lemma of Fox [23] which implies $\mathsf{RS}_{n}(cn)\leq n/\log^{(\ell)}$ for $\ell=O(\log(1/c))$ where $\log^{(x)}$ is the iterated log function. We note that this result is implicit in [23] and was mentioned in the paper of [24]. To our knowledge, this upper bound does not carry over to ORS graphs (we briefly discuss this at the beginning of Section 5). Our second result is a similar upper bound for ORS albeit with a worse dependence on constant $c$.
@@
Since every RS graph is also an ORS graph with the same parameters, Theorem 2 immediately implies the same upper bound for $\mathsf{RS}_{n}(cn)$. Note that this implication is not a new result, but the proof is very different from the abovementioned upper bound.
@@ 1.2 Perspective: ORS vs RS
Does it hold that $\mathsf{ORS}_{n}(cn)=\operatorname{poly}(\mathsf{RS}_{n}(O(cn)))$?
@@
In the event that the answer to 2 is negative, one might wonder whether we can improve Theorem 1 by parametrizing it based on $\mathsf{RS}$ instead of $\mathsf{ORS}$. Put differently, suppose that $\mathsf{ORS}_{n}(cn)=n^{1-o(1)}$ and $\mathsf{RS}_{n}(cn)=n^{o(1)}$. Can we somehow utilize the sparsity of RS graphs (instead of ORS graphs) in this case to improve existing dynamic matching algorithms? We start Section 2 by providing an input construction which informally shows ORS is the right parameter for Theorem 1 even if RS is much sparser.
@@ 2 Our Techniques
Before describing the intuition behind our algorithm of Theorem 1, let us start with a sequence of updates that, in a sense, explains why existence of dense ORS graphs would make it challenging to maintain a $(1-\varepsilon)$-approximate maximum matching in a fully dynamic setting. We then proceed to show that this input construction is essentially the only instance that we do not know how to solve fast.
An input construction based on ORS graphs:
@@
The input construction above implies that to maintain a $(1-\varepsilon)$-approximation of maximum matching, either we have to find a way to identify induced matchings of an ORS graph fast (without scanning the neighbors of each vertex) or have to parameterize our algorithm’s update-time by $t$, the density of ORS graphs.
@@
Let us for this informal overview of our algorithm assume that the maximum matching size is at least $\Omega(n)$. Having this assumption allows us to find the matching once, do nothing for the next $\varepsilon n$ updates, and then repeat without hurting the size of the approximate matching that we find by more than a $1+O(\varepsilon)$ factor.
@@
For the first problem discussed above, we present an algorithm that runs in (essentially) $O(n^{2}/d)$ time to find the maximum matching in $G[U]$. Here $d$ is a parameter that depends on the structure of $G[U]$ that measures how easy it is to find an approximate maximum matching of $G[U]$. Intuitively, if the average degree within $G[U]$ is $d$, we can random sample pairs of vertices to add to the matching. If each vertex is adjacent to $d$ others, we only need $O(n/d)$ samples to match it and the algorithm runs in $O(n^{2}/d)$ time. Of course, this can take up to $\Omega(n^{2})$ time if $G[U]$ is sparse – e.g. when it is an induced matching. Then instead of charging a matching in $G[U]$ that is an induced matching, we charge this matching of average degree at most $d$ inside. Let $M_{1},\ldots,M_{t}$ be the matchings that we charge and let $d_{i}$ be the average degree of the $i$-th matching and suppose that these matchings are edge-disjoint. We show in our update-time analysis that $\sum_{i=1}^{t}1/d_{i}$ can be at most $\mathsf{ORS}(\varepsilon^{2}n)$, and therefore the total time spent by our algorithm can be upper bounded by $n^{2}\mathsf{ORS}(\varepsilon^{2}n)$.
For the second problem, or in other words, to ensure that the matchings that we charge are edge-disjoint, we maintain a set $S$ and add all edges of any matching that we charge to this set. Thereafter, before solving $G[U]$, we first go over the edges stored in this set and see whether they can be used to find a large matching in $G[U]$. If they do, we do not run the random sampling algorithm discussed above. If not, the matching that we find must be edge-disjoint. We have to be careful that we do not make $S$ too dense though as we spend linear time in the size of $S$. Our final algorithm resets $S$ after a certain number of updates.
@@
To obtain our upper bound of Theorem 2, we partition the matchings into two subsets, $\mathcal{M}$ and $\mathcal{M}^{\prime}$, based on their order in the sequence. A key insight is the following: take a vertex $v$ and suppose that it is matched by some matching $M$ in $\mathcal{M}^{\prime}$. If we remove all neighbors of vertex $v$ in $\mathcal{M}$, then it can be proved that no vertex of matching $M$ is removed because otherwise there must be an edge from $\mathcal{M}$ that matches two vertices of $M$, violating the inducedness property. Intuitively, this shows that if vertices have large degrees in $\mathcal{M}$, we can remove a relatively large number of vertices without hurting the matchings that include this vertex. To derive the upper bound, we carefully select a set of pivots based on the degrees in $\mathcal{M}$, and remove the neighbors of these pivots. We show that this reduces the number of vertices significantly enough, and keeps the size of a small (but sufficiently many) of the matchings unchanged. If the initial number of matchings is so large, we show that we can iteratively applying this procedure. Because the size of matchings do not change but the number of vertices drops, we get that the process should eventually stop. This implies the upper bound on the number of matchings in the starting graph.

### v3

@@ Abstract
We study the fully dynamic maximum matching problem. In this problem, the goal is to efficiently maintain an approximate maximum matching of a graph that is subject to edge insertions and deletions. Our focus is particularly on algorithms that maintain the edges of a $(1-\varepsilon)$-approximate maximum matching for an arbitrarily small constant $\varepsilon>0$. Until recently, the fastest known algorithm for this problem required $\Theta(n)$ time per update where $n$ is the number of vertices. This bound was slightly improved to $n/(\log^{*}n)^{\Omega(1)}$ by Assadi, Behnezhad, Khanna, and Li [STOC’23] and very recently to $n/2^{\Omega(\sqrt{\log n})}$ by Liu [FOCS’24]. Whether this can be improved to $n^{1-\Omega(1)}$ remains a major open problem.
In this paper, we introduce Ordered Ruzsa-Szemerédi (ORS) graphs (a generalization of Ruzsa-Szemerédi graphs) and show that the complexity of dynamic matching is closely tied to them. For $\delta>0$, define $\mathsf{ORS}(\delta n)$ to be the maximum number of matchings $M_{1},\ldots,M_{t}$, each of size $\delta n$, that one can pack in an $n$-vertex graph such that each matching $M_{i}$ is an induced matching in subgraph $M_{1}\cup\ldots\cup M_{i}$. We show that there is a randomized algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching of a fully dynamic graph in
$\widetilde{O}\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}\right)$
amortized update-time.
While the value of $\mathsf{ORS}(\Theta(n))$ remains unknown and is only upper bounded by $n^{1-o(1)}$, the densest construction known from more than two decades ago only achieves $\mathsf{ORS}(\Theta(n))\geq n^{1/\Theta(\log\log n)}=n^{o(1)}$ [Fischer et al. STOC’02]. If this is close to the right bound, then our algorithm achieves an update-time of $\sqrt{n^{1+O(\varepsilon)}}$, resolving the aforementioned longstanding open problem in dynamic algorithms in a strong sense.
@@ 1 Introduction
The dynamic matching problem has received a lot of attention over the last two decades [39, 8, 9, 38, 32, 23, 16, 17, 22, 21, 41, 26, 2, 18, 13, 15, 43, 19, 20, 40, 34, 31, 12, 25, 11, 7, 24]. There is a relatively simple algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching, for any fixed $\varepsilon>0$, in just $O(n)$ time per update (see e.g. [32]). The update-time can be significantly improved if we worsen the approximation. For instance, a 1/2-approximation can be maintained in $\operatorname{poly}\log n$ time per update [42, 8, 13], or an (almost) 2/3-approximation can be maintained in $O(\sqrt{n})$ time [17]. Nonetheless, when it comes to algorithms with approximation ratio better than $2/3$, the update-time stays close to $n$. A slightly sublinear algorithm was proposed by Assadi, Behnezhad, Khanna, and Li [6] which runs in $n/(\log^{*}n)^{\Omega(1)}$ time per update and maintains a $(1-o(1))$-approximation. In a very recent paper, Liu [36] improved this to $n/2^{\Omega(\sqrt{\log n})}$ via a nice connection to algorithms for the online matrix-vector multiplication (OMv) problem. Despite this progress, the following remains a major open problem:
@@
Is it possible to maintain a $(1-\varepsilon)$-approximate maximum matching in a fully dynamic graph, for any fixed $\varepsilon>0$, in $n^{1-\Omega(1)}$ update-time?
We note that there is an orthogonal line of work on fully dynamic algorithms that instead of maintaining the edges of the matching, maintains only its size [11, 25, 24]. For this easier version of the problem, Bhattacharya, Kiss, and Saranurak [24] positively resolved the open problem above. However, their algorithm crucially relies on only estimating the size and does not work for the problem of maintaining the edges of the matching. We refer interested readers to [11] where the difference between the two versions of the problem is mentioned.
@@ 1.1 Our Contributions
Let $\varepsilon>0$ be fixed. There is a fully dynamic algorithm that maintains the edges of a $(1-\varepsilon)$-approximate maximum matching in $O\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}\operatorname{poly}(\log n)\right)$ amortized update-time. The algorithm is randomized but works against adaptive adversaries.
@@
The best lower bound on $\mathsf{RS}_{n}(cn)$— i.e., the densest known construction—is that of [27] from more than two decades ago which shows $\mathsf{RS}_{n}(cn)\geq n^{\Omega_{c}(1/\log\log n)}=n^{o(1)}$. This is indeed the densest known construction of ORS graphs we are aware of too. If this turns out to be the right bound, Theorem 1 implies a $(1-\varepsilon)$-approximation in $\widetilde{O}(n^{1/2+\varepsilon})$ time, an almost quadratic improvement over prior near-linear in $n$ algorithms of [6, 36]. In fact, we note that so long as $\mathsf{ORS}_{n}(cn)$ is moderately smaller than $n$ (say $\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))\ll n^{1-\Omega(1)}$) Theorem 1 still implies a truly sublinear in $n$ update-time algorithm, positively resolving 1.
Finally, we note that there is a long body of work on proving conditional lower bounds for dynamic problems. For instance, the OMv conjecture can be used to prove that maintaining an exact maximum matching requires near-linear in $n$ update-time [33]. Adapting these lower bounds to the $(1-\varepsilon)$-approximate maximum matching problem has remained open since then. Our Theorem 1 implies that proving such lower bounds either requires a strong lower bound of near-linear on $\mathsf{ORS}_{n}(\Theta(n))$, or requires a conjecture that implies this.
Contribution 2: Upper Bounding ORS.
Unfortunately there is a huge gap between existing lower and upper bounds for $\mathsf{RS}_{n}(cn)$ (and as a result also for $\mathsf{ORS}_{n}(cn)$). The best known upper bound on $\mathsf{RS}_{n}(cn)$ for linear size matchings follows from the improved triangle-removal lemma of Fox [28] which implies $\mathsf{RS}_{n}(cn)\leq n/\log^{(\ell)}n$ for $\ell=O(\log(1/c))$ where $\log^{(x)}$ is the iterated log function. We note that this result is implicit in [28] and was mentioned in the paper of [29]. To our knowledge, this upper bound does not carry over to ORS graphs (we briefly discuss this at the beginning of Section 6). Our second result is a similar upper bound for ORS albeit with a worse dependence on constant $c$.
@@
Since every RS graph is also an ORS graph with the same parameters, Theorem 2 immediately implies the same upper bound for $\mathsf{RS}_{n}(cn)$. Note that this implication is not a new result, but the proof is different from that of [28] and is closer to the arguments in [29].
@@ 1.2 Perspective: ORS vs RS
Is it true that for fixed $\varepsilon>0$,
$\mathsf{ORS}_{n}(\varepsilon n)\leq\operatorname{poly}(\mathsf{RS}_{n}(\Theta_{\varepsilon}(n)))?$
@@
In the event that the answer to 2 is negative, one might wonder whether we can improve Theorem 1 by parametrizing it based on RS instead of ORS. Put differently, suppose that $\mathsf{ORS}_{n}(cn)=n^{1-o(1)}$ and $\mathsf{RS}_{n}(cn)=n^{o(1)}$. Can we somehow utilize the sparsity of RS graphs (instead of ORS graphs) in this case to improve existing dynamic matching algorithms? We start Section 2 by providing an input construction which informally shows ORS is the right parameter for Theorem 1 even if RS graphs turn out to be much sparser.
1.3 Connections and Recent Developments
Connection to Sublinear Time Algorithms:
Our algorithm crucially relies on sublinear time algorithms; particularly the algorithm of the first author in [10] for estimating the size of maximum matching. While sublinear-time algorithms have also played a crucial role in the dynamic matching algorithms of [11, 25, 24], these algorithms maintain only the size of the matching as opposed to its edges. Our algorithm, on the other hand, maintains the matching explicitly yet relies on matching size estimators crucially. On a high level, our algorithm can tolerate the time needed to find a large matching in a specific induced subgraph so long as a large matching is guaranteed to exist there. We use the fast sublinear-time matching algorithm of [10] to first verify existence of a large matching, then spend the time needed if this matching is large.
Connection to a Lower Bound of Liu [36]:
In his very recent work, Liu [36] showed that under the approximate OMv conjecture, there is no algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching in $n^{1-\Omega(1)}\operatorname{poly}(1/\varepsilon)$ update-time. This might, at the first glance, appear to contradict Theorem 1 which runs in $n^{1-\Omega(1)}$ time if it so happens that $\mathsf{ORS}(\Theta(n))=n^{1-\Omega(1)}$. However, we note that there is no contradiction even if approximate OMv conjecture turns out to be true and $\mathsf{ORS}(\Theta(n))=n^{1-\Omega(1)}$ at the same time! The reason is that when parameter $\varepsilon$ is sufficiently sub-constant, dense constructions of RS (and consequently ORS) with $\Theta(n^{2})$ edges already exist [1]. So our Theorem 1 is only potentially useful for fixed (or mildly sub-constant) $\varepsilon$. On the other hand, the reduction of [36] sets $\varepsilon=n^{-\Omega(1)}$, thus only targets the sub-constant regime.
Interestingly, while the result of [36] removes the combinatorial structure of the dynamic matching problem and reduces it to a purely algebraic question for the sub-constant regime of $\varepsilon$, our work shows that the fixed regime of $\varepsilon$ is, in fact, a completely combinatorial question.
Recent Developments:
After the first version of this paper appeared online, the nice follow-up works of Assadi and Khanna [4] and Kiss [35] improved our update-time from $O(\sqrt{n^{1+o(1)}\mathsf{ORS}(\Theta(n)})$ to $O(n^{o(1)}\mathsf{ORS}(\Theta(n)))$. Importantly, their algorithms run in $n^{o(1)}$ update-time under $\mathsf{ORS}(\Theta(n))=n^{o(1)}$, showing that ORS graphs fully characterize the hardness of the dynamic matching problem.
@@ 2 Our Techniques
In this section, we provide an informal overview of our algorithm for Theorem 1 as well as the upper bound of Theorem 2.
Before describing the intuition behind our algorithm of Theorem 1, let us start with a sequence of updates that, in a sense, explains why existence of dense ORS graphs would make it challenging to maintain a $(1-\varepsilon)$-approximate maximum matching in a fully dynamic setting.
Why ORS graphs are seemingly hard:
@@
The input construction above implies that to maintain a $(1-\varepsilon)$-approximation of maximum matching, either we have to find a way to identify induced matchings of an ORS graph fast (without scanning the neighbors of each vertex) which appears extremely challenging, or we have to parameterize our algorithm’s update-time by $t$, the density of ORS graphs. We take the latter approach in this work.
@@
Let us for this informal overview of our algorithm assume that the maximum matching size is at least $\Omega(n)$. Having this assumption (which comes w.l.o.g. as stated by Proposition 5) allows us to find the matching once, do nothing for the next $\varepsilon n$ updates, and then repeat without hurting the size of the approximate matching that we find by more than a $1+O(\varepsilon)$ factor.
@@
For the first problem discussed above, we present an algorithm that runs in (essentially) $O(n^{2}/d)$ time to find the maximum matching in $G[U]$. Here $d$ is a parameter that depends on the structure of $G[U]$ that measures how easy it is to find an approximate maximum matching of $G[U]$. Intuitively, if the average degree within $G[U]$ is $d$, we can random sample pairs of vertices to add to the matching. If each vertex is adjacent to $d$ others, we only need $O(n/d)$ samples to match it and the algorithm runs in $O(n^{2}/d)$ time. Of course, this can take up to $\Omega(n^{2})$ time if $G[U]$ is sparse – e.g. when it is an induced matching. Then instead of charging a matching in $G[U]$ that is an induced matching, we charge this matching of average degree at most $d$ inside, which we call a certificate. Let $M_{1},\ldots,M_{t}$ be the certificates that we charge and let $d_{i}$ be the average degree of the $i$-th matching and suppose that these matchings are edge-disjoint. We show in our update-time analysis that $\sum_{i=1}^{t}1/d_{i}$ can be at most $\mathsf{ORS}(\Theta(n))$, and therefore the total time spent by our algorithm during a phase can be upper bounded by $n^{2}\mathsf{ORS}(\Theta(n))$.
For the second problem, or in other words, to ensure that the certificate matchings that we charge are edge-disjoint, we maintain a set $H_{\text{cert}}$ and add all edges of any matching that we charge to this set. Thereafter, before solving $G[U]$, we first go over the edges stored in this certificate set and see whether they can be used to find a large matching in $G[U]$. If they do, we do not run the random sampling algorithm discussed above. If not, the matching that we find must be edge-disjoint. We have to be careful that we do not make $H_{\text{cert}}$ too dense though as we spend linear time in the size of $H_{\text{cert}}$. Our final algorithm resets $H_{\text{cert}}$ after a certain number of updates.
We note that our informal discussion of this section hides many (important) details of the final algorithm that we formalize in Section 4.2.
@@
To obtain our upper bound of Theorem 2, we partition the matchings into two subsets, $\mathcal{M}$ and $\mathcal{M}^{\prime}$, based on their order in the sequence. A key insight (a variant of which was used in the RS upper bound of [29]) is the following: take a vertex $v$ and suppose that it is matched by some matching $M$ in $\mathcal{M}^{\prime}$. If we remove all neighbors of vertex $v$ in $\mathcal{M}$, then it can be proved that no vertex of matching $M$ is removed because otherwise there must be an edge from $\mathcal{M}$ that matches two vertices of $M$, violating the inducedness property. Intuitively, this shows that if vertices have large degrees in $\mathcal{M}$, we can remove a relatively large number of vertices without hurting the matchings that include this vertex. To derive the upper bound, we carefully select a set of pivots based on the degrees in $\mathcal{M}$, and remove the neighbors of these pivots. We show that this reduces the number of vertices significantly enough, and keeps the size of a small (but sufficiently many) of the matchings unchanged. If the initial number of matchings is so large, we show that we can iteratively applying this procedure. Because the size of matchings do not change but the number of vertices drops, we get that the process should eventually stop. This implies the upper bound on the number of matchings in the starting graph.